OpenStudy (anonymous):

Prove: 3 divides (n^(3) - n) n is an element of N

OpenStudy (anonymous):

you can use Fermat's Little Theorem n^(p-1) = 1 (mod p) so n^p = a (mod p) n^3 = n (mod 3) so 3|n^3 - n or doing it the other way factor the n^3 - n = n(n^2 - 1) = n(n+1)(n-1) if 3 divides n^3 - n then it divides n(n+1)(n-1) so this means at least one of these n, n+1, n-1 is divisible by 3 and you just need to check it at 1,2,3 when n = 1 you get 1, 2, 0 = 0 divisible by 3 n= 2 you get 2, 3, 1 = 3 divisble by 3 n = 3 you get 3, 4, 2 = 3 divislbe by 3 and the other method i would use is induction base case: n= 1 so 1^3 - 1 = 0 true the assume arbitrary k where k^3 - k = 3x where x in Z and prove for k+1 (k+1)^3 - (k+1) = 3x (k+1)((k+1)^2 - 1) (k+1)(k^2 + 2k + 1 - 1) (k+1)(k^2 + 2k) = 3x k^3 + 2k^2 + k^2 + 2k = k^3 + 3k^2 + 2k = k^3 - k + 3k^2 + 3k = 3x by inductive hypothesis 3z + 3k^2 + 3k = 3x where z is an integer so 3(z + k^2 + k) = 3x so 3 divides this equation so this holds for k+1 which holds for n

OpenStudy (anonymous):

i might be doing the 2nd method wrong so i would recommend induction instead

OpenStudy (anonymous):

Can you show me induction with this problem?

OpenStudy (anonymous):

i showed it as the 3rd method you want to do base case: n= 1 so 1^3 - 1 = 0 true the assume arbitrary k where \[3 | k^3 - k\] and prove for k+1 \[3|(k+1)^3 - (k+1)\]

OpenStudy (anonymous):

divisor also means \[(k+1)^3 - (k+1) = 3x\] where \[x \in \mathbb{Z}\] so first factor out a k+1 \[(k+1)((k+1)^2 - 1) = 3x\]

OpenStudy (anonymous):

So what step did you think was wrong. ^

OpenStudy (anonymous):

then multiple the inside out \[(k+1)(k^2 + 2k + 1 - 1) = (k+1)(k^2 + 2k) = 3x\] distribute again \[k^3 + 2k^2 + k^2 + 2k = k^3 + 3k^2 + 2k = 3x\]

OpenStudy (anonymous):

The induction is did in the first post is fine its just the 2nd method that i think it was wrong im showing you step by step on how to do this with induction, what part are you confused about?

OpenStudy (anonymous):

I am confused about what you said is wrong. I do not want to learn it incorrect.

OpenStudy (anonymous):

anyways you now have \[k^3 + 3k^2 + 2k = 3x\] add and subtract k on the LHS \[k^3 + 3k^2 + 2k + k - k= 3x\] \[(k^3 - k)+ 3k^2 + 3k= 3x\] the by inductive hypothesis 3z + 3k^2 + 3k = 3x where z is an integer factor out a 3 3(z+k^2 + k) = 3x so this holds for k+1 which implies it holds for n

OpenStudy (anonymous):

i showed you all the correct steps for the induction problem i hope you are reading what i am writing/understanding just look at my 3rd post and start from there

OpenStudy (anonymous):

I understand.

OpenStudy (anonymous):

alright then glad to help